(1 point) What nominal rate compounded continuously will increase your investment b.25% over 8 years? Answer:
Answer #1

Answer:

FV = P * e rt (P = Principal, r = Rate of Interest and t = Time period)

Let us assume initial investment = 1

Hence:

FV = 1 * (1 + 25%) = 1.25

Hence:

1.25 = e r * 8

since e = 2.718281828459

=>8r = 0.223

=> r = 0.223 / 8 =2.79%

2.79 %